OpenStudy (anonymous):

Sara had 7/8 pounds of turkey. She used 2/3 pound for sandwiches. How many pounds of turkey does Sara have left?

OpenStudy (inowalst):

Make the denominators the same. \[\frac{ 7 }{ 8 }\times \frac{ 3 }{ 3 }\] Then: \[\frac{ 2 }{ 3 }\times \frac{ 8 }{ 8 }\] Can you multiply those fractions and tell me what you get for each?

OpenStudy (anonymous):

not really. im not completely understanding why you would put 8 over 8 and 3 over 3

OpenStudy (inowalst):

Oh because we wanna get the denominators to be the same, in order to subtract the fractions.

OpenStudy (anonymous):

so........ \[21 \over 24\] and \[16 \over 24\] ?

OpenStudy (inowalst):

Yes, now subtract.

OpenStudy (anonymous):

\[5 \over 24?\]

OpenStudy (inowalst):

Yes. That's your answer.

OpenStudy (anonymous):

thx :)

OpenStudy (inowalst):

You're welcome!